# Generalization Below the Edge of Stability: The Role of Data Geometry

### v1

@@ 4.3 Empirical Evidence for the Data Shatterability Principle
This experimental result validates not only our main theorems but also the core techniques of our proof. To illustrate this, we introduce Tukey depth $\mathrm{depth}_{\mathcal{P}_{{\bm{X}}}}({\bm{x}}):=\inf_{\|{\bm{u}}\|_{2}=1}\mathbb{P}\big({\bm{u}}^{\top}{\bm{X}}\geq{\bm{u}}^{\mathsf{T}}{\bm{x}})$, which measures the centrality of a point $x$ by finding the minimum data mass (either population or empirical) on one side of any hyperplane passing through it (Tukey, 1975). Our key claim is that deeper regions of a distribution are hard-to-shatter. For a ReLU ridge to introduce non-linearity, or “wiggleness”, within a deep region $\Omega_{T}:=\{{\bm{x}}:\mathrm{depth}_{\mathcal{P}_{{\bm{X}}}}({\bm{x}})\geq T\}$, its decision boundary must pass through that region. By the very definition of Tukey depth, the corresponding neuron is then guaranteed to activate on at least a $T$-fraction of the data. This provides a lower bound on the EoS weight function $g$ for these specific neurons that contribute to the function’s nonlinearity in the $\Omega_{T}$.
Consequently, within this deep region, the stability-induced weighted path norm constraint effectively becomes a more traditional unweighted path norm bound for the part of the network creating local complexity. Such function classes are known to ensure generalization (Parhi and Nowak, 2023; Neyshabur et al., 2015). Outside this core, the BEoS constraint provides no meaningful control for neurons activating there, as $g$ can be vanishingly small, allowing for large unweighted norms that facilitate memorization. Heuristically, deeper regions should generalize better. The right panel of Figure 4 provides a striking visual confirmation of this principle on a real-world dataset.
Figure 4: Data geometry and memorization on MNIST. Left panel: Comparison of training curves under the same ground-truth predictor with Gaussian inputs versus MNIST inputs ($n=30000$). GD on the Gaussian data set quickly interpolates, while MNIST resists overfitting for tens of thousands of steps. Right panel: Prediction error against Tukey depth for MNIST samples. Shallow points (low depth) exhibit larger errors. This region refers to “highly shatterable region”.
Our upper-bound proof technique, which refines the domain decomposition strategy from (Liang et al., 2025), operationalizes this split. We partition the data space into a deep core $\Omega_{T}$ and its shallow complement. We control the generalization error in the core, while conservatively bounding the error in the shallow region by its total probability mass. The final bound arises from optimizing the trade-off in selecting $T$. Crucially, this strategy is consistent with our lower-bound construction, as well as the "neural shattering" phenomenon analyzed in (Liang et al., 2025). The “hard-to-learn” functions in Construction H.4 are built with neurons that activate exclusively in the low-depth regions. This shows that surrendering the shallow region in the upper bound is not a mere mathematical trick, but reflects a fundamental characteristic of the BEoS regime. In the extreme case of data on a sphere, where all points have zero depth, this technique becomes inapplicable—which aligns perfectly with our finding that stable interpolation is possible in that setting.
@@ 5 Discussion and Further Questions
In this work, we present a mechanism explaining how data geometry governs the implicit bias of neural networks trained below the Edge of Stability. We introduce the principle of “data shatterability,” demonstrating that geometries resistant to shattering guide gradient descent towards discovering shared, generalizable representations. Conversely, we show that easily shattered geometries, such as data concentrated on a sphere, permit stable solutions that memorize the training data.
Our framework opens several promising avenues for future research. A central question is the connection between shatterability and optimization. The observation that a flip side of being prone to overfitting is often faster optimization leads to a natural hypothesis: are high-shatterability distributions easier to optimize? This, in turn, raises further questions about the role of normalization techniques. For instance, do normalization techniques like Batch Norm accelerate training precisely by enforcing more isotropic, and thus more shatterable, representations at each layer? This line of inquiry extends naturally to deep networks, where hidden layers not only sense the initial data geometry but actively create a new “representation geometry”. Can our principles be translated to the understanding of representation geometry? Finally, this framework may offer a new lens to understand architectural inductive biases. For example, do CNNs generalize well precisely because their local receptive fields impose an architectural constraint that inherently reduces the model’s ability to shatter the data, forcing it to learn local, reusable features? Answering such questions, alongside developing a quantifiable metric for shatterability, remains a key direction.
@@ Appendix A The Use of Large Language Models (LLMs)
Appendix B More Related Works
@@
Edge of Stability (EoS) and minima stability. The EoS literature primarily seeks to explain when and why training operates near instability and how optimization proceeds there [Cohen et al., 2020, Kong and Tao, 2020, Arora et al., 2022, Ahn et al., 2022, Damian et al., 2024]. Central flows offer an alternative viewpoint on optimization trajectories that also emphasizes near-instability behavior [Cohen et al., 2025]. Closest to our work is the line on minima stability [Ma and Ying, 2021, Mulayoff et al., 2021, Nacson et al., 2023, Wu and Su, 2023, Qiao et al., 2024], which links Hessian spectra and training noise to the geometry of solutions but largely leaves generalization out of scope. We leverage the EoS/minima-stability phenomena to define and analyze a data-distribution-aware notion of stability, showing adaptivity to low-dimensional structure and making explicit how distributional geometry shapes which stable minima GD selects.
Flatness vs. generalization. Whether (and which notion of) flatness predicts generalization remains debated. Several works argue sharp minima can still generalize [Dinh et al., 2017], propose information-geometric or Fisher–Rao–based notions [Liang et al., 2019], or develop relative/scale-invariant flatness measures [Petzka et al., 2021]. We focus on the largest curvature direction (i.e., $\lambda_{\max}$) motivated by EoS/minima-stability. Our results rigorously prove that flatness in this notion does imply generalization (note that there is no contradiction with Dinh et al. [2017]), but it depends on data distribution.
@@ Appendix C Details of Experiments
(b) Sample-wise error versus Tukey depth at epoch 5,0005{,}000
@@ D.2 Total Variation Semi-norm on Radon Domain
We now connect the (unweighted) variation semi-norm of shallow ReLU networks to an analytic description on the Radon domain. Our presentation follows [Parhi and Nowak, 2021, 2023].
@@
The Radon framework encodes a function $f$ by its integrals over affine hyperplanes faithfully in the senses that the Radon transform is invertible up to a known dimension-dependent constant via a one-dimensional “ramp” filter in $t$.
@@
$c_{d}\,f\;=\;\mathscr{R}^{*}\left\{\,\Lambda_{d-1}\mathscr{R}f\,\right\},$
where $\Lambda_{d-1}$ acts in the $t$-variable with Fourier symbol $\widehat{\Lambda_{d-1}\Phi}({\bm{u}},\omega)=i^{\,d-1}\,|\omega|^{\,d-1}\,\hat{\Phi}({\bm{u}},\omega)$.
The inversion formula motivates measuring the “ridge-curvature” of $f$ by differentiating in the Radon offset $t$ after filtering, and aggregating its magnitude over all orientations and offsets.
The next definition is the sole norm we need on the Radon domain; it specializes all higher-order variants to the ReLU case.
@@
where the fractional power is understood in the tempered-distribution sense. The null space of $\mathscr{R}\mathrm{TV}^{2}(\cdot)$ is the set of affine functions on $\mathbb{R}^{d}$.
Proposition D.7 (Equivalence of seminorms on bounded domains [Parhi and Nowak, 2021]).
@@
The key structural reason is simple: $\partial_{t}^{2}\Lambda_{d-1}\mathscr{R}$ turns each ReLU ridge $\phi({\bm{u}}^{\mathsf{T}}{\bm{x}}-t)$ into a Dirac mass at $({\bm{u}},t)$ on $\mathbb{S}^{d-1}\times\mathbb{R}$}, so superpositions of ridges correspond exactly to finite signed measures on $\mathbb{S}^{d-1}\times\mathbb{R}$}, and the total-variation of that measure coincides with both the variation seminorm and $\mathscr{R}\mathrm{TV}^{2}(\cdot)$ after fixing the affine null space.
Remark D.8 (Takeaway).
For ReLU networks on bounded domains, the three viewpoints
$\text{path-norm }\left\lVert f\right\rVert_{\mathrm{path}}\;\;\longleftrightarrow\;\;\text{unweighted variation }|f|_{\mathrm{V}}\;\;\longleftrightarrow\;\;\text{Radon-TV }\mathscr{R}\mathrm{TV}(f)$
are equivalent up to the affine null space. We will freely switch between them in the sequel.
@@ E.2 Empirical Process for the Weight Function $g$
Taking the supremum over $({\bm{u}},t)\in\mathbb{S}^{d-1}\times[-1,1]$ and applying Lemmas E.4 and E.5 with a union bound yields
@@
Finally, the equivalence $g\asymp p\,s$ transfers this bound to $\big|g_{\mathcal{D}}-g_{\mathcal{P}}\big|$ at the cost of an absolute multiplicative factor and one more failure event. ∎
@@ Appendix F Generalization Upper Bound: Mixture of Low-Dimensional Balls
In this section, we present the proof of Theorem 3.2. First, we prove the simple case of singe-subspace assumption ($J=1$) via Theorem F.2.
@@ F.1 Case: uniform distribution on unit disc of a linear subspace
By Lemma F.1, it remains to consider the case of $\operatorname{proj}_{V}^{*}f_{{\bm{\theta}}}$. Similarly, for any ${\bm{u}}\in\mathbb{S}^{d-1}$ and any data set $\mathcal{D}\subset V$, we have $g({\bm{u}},t)=g(\operatorname{proj}_{V}({\bm{u}}),t)$. Therefore, we just need to consider the generalization gap with respect to the ${\bm{\Theta}}^{V}_{g_{\mathcal{D}}}(\mathbb{B}_{1}^{V_{j}};M,C)=\left\{\operatorname{proj}_{V}^{*}f_{{\bm{\theta}}}\>\colon\>f_{{\bm{\theta}}}\in{\bm{\Theta}}_{g_{\mathcal{D}}}(\mathbb{B}_{1}^{V_{j}};M,C)\right\}$. Therefore, we just need consider the case where the whole algorithm with any dataset sample from $V$ operates in $V$ and we get the result from [Liang et al., 2025, Theorem F.8] by replacing $\mathbb{R}^{d}$ with $V\cong\mathbb{R}^{m}$. ∎

### v2

@@ 5 Discussion and Further Questions
In this work, we present a mechanism explaining how data geometry governs the implicit bias of neural networks trained below the Edge of Stability. We introduce the principle of “data shatterability,” demonstrating that geometries resistant to shattering guide gradient descent towards discovering shared, generalizable representations.
Limitations. While our theoretical framework instantiates the principle of data shatterability, there are limitations to our current analysis. First, the proposed half-space-depth concentration index ($\mathsf{S}_{\text{DQ}}$) serves as a proxy primarily for isotropic distributions and extending this scalar metric to quantify the shatterability of arbitrary, anisotropic, structured data remains a non-trivial challenge. Second, our results are derived for two-layer ReLU networks to maintain tractability in the function-space analysis. Generalizing these bounds to deep and complicated networks presents substantial theoretical hurdles, particularly in characterizing how the EoS-induced regularity constraint propagates through hierarchical layers without becoming ensuring the bounds remain non-vacuous.
Further Questions. Our framework opens several promising avenues for future research. A central question is the connection between shatterability and optimization. The observation that a flip side of being prone to overfitting is often faster optimization leads to a natural hypothesis: are high-shatterability distributions easier to optimize? This, in turn, raises further questions about the role of normalization techniques. For instance, do normalization techniques like Batch Norm accelerate training precisely by enforcing more isotropic, and thus more shatterable, representations at each layer? This line of inquiry extends naturally to deep networks, where hidden layers not only sense the initial data geometry but actively create a new “representation geometry”. Can our principles be translated to the understanding of representation geometry? Finally, this framework may offer a new lens to understand architectural inductive biases. For example, do CNNs generalize well precisely because their local receptive fields impose an architectural constraint that inherently reduces the model’s ability to shatter the data, forcing it to learn local, reusable features? Answering such questions, alongside developing a quantifiable metric for shatterability, remains a key direction.
6 Acknowledgments
The research was partially supported by NSF Award # 2134214. Tongtong Liang thanks Dan Qiao and Zihan Shao for providing helpful suggestions.
@@ Appendix A The Use of Large Language Models (LLMs)
Appendix B More Related Works and Discussions
B.1 More Related Works
@@
Edge of Stability (EoS) and minima stability. The EoS literature primarily seeks to explain when and why training operates near instability and how optimization proceeds there [Cohen et al., 2020, Kong and Tao, 2020, Arora et al., 2022, Ahn et al., 2022, Damian et al., 2024]. Central flows offer an alternative viewpoint on optimization trajectories that also emphasizes near-instability behavior [Cohen et al., 2025]. Closest to our work is the line on minima stability [Ma and Ying, 2021, Mulayoff et al., 2021, Nacson et al., 2023], which links Hessian spectra and training noise to the geometry of solutions but largely leaves generalization out of scope. We leverage the EoS/minima-stability phenomena to define and analyze a data-distribution-aware notion of stability, showing adaptivity to low-dimensional structure and making explicit how distributional geometry shapes which stable minima GD selects.
Nonparametric function estimation with neural networks.
The notion of generalization gap is closely related to the estimation error. It is well known that neural networks are minimax optimal estimators for a wide variety of functions [Suzuki, 2018, Schmidt-Hieber, 2020, Kohler and Langer, 2021, Parhi and Nowak, 2023, Zhang and Wang, 2023, Wu and Su, 2023, Yang and Zhou, 2024, Qiao et al., 2024]. Outside of the univariate work of Qiao et al. [2024], all prior works construct their estimators via empirical risk minimization problems. Thus, they do not incorporate the training dynamics that arise when training neural networks in practice. In contrast, Qiao et al. [2024], Liang et al. [2025] derive nonparametric guarantees for solutions selected directly by gradient descent dynamics, without assuming even stationary condition. Our work further develops in this direction by modeling how data geometry affect generalization for the gradient trajectories below the edga-of-stability.
Flatness vs. generalization. Whether (and which notion of) flatness predicts generalization remains debated. Several works argue sharp minima can still generalize [Dinh et al., 2017], propose information-geometric or Fisher-Rao-based notions [Liang et al., 2019], or develop relative/scale-invariant flatness measures [Petzka et al., 2021]. We focus on the largest curvature direction (i.e., $\lambda_{\max}$) motivated by EoS/minima-stability. Our results rigorously prove that flatness in this notion does imply generalization (note that there is no contradiction with Dinh et al. [2017]), but it depends on data distribution.
@@
B.2 A Gradient-dynamics Perspective on Stability and Data Geometry
We now discuss our results from the viewpoint of gradient dynamics.
Why Shattering Neural Networks May Be Dynamically Stable. Here we paraphrase and adapt the discussion from [Liang et al., 2025, Appendix A.4]. The BEoS condition defines a set of dynamically stable parameter states, and the data-dependent weight function $g_{\mathcal{D}}({\bm{u}},t)$ provides a static summary of how expensive it is to place a ReLU ridge at orientation $\bm{u}$} and threshold $t$. Small values of $g_{\mathcal{D}}({\bm{u}},t)$ indicate weak stability constraints in that region of parameter space, so a neuron aligned with $({\bm{u}},t)$ can carry a large coefficient while still satisfying the BEoS curvature bound. In highly shatterable geometries, the shallow shell contains many such directions with tiny $g_{\mathcal{D}}$, creating ample room inside the stable set for high-magnitude, sparsely activating neurons supported on disjoint caps.
This static picture is closely tied to the actual gradient dynamics. If the dataset is highly shatterable in the sense of the half-space-depth concentration index, a neuron’s activation boundary can easily drift toward regions where it fires on only a few data points. Once those few points are already well-fitted, the gradient contributions in (13) become small and localized, so the neuron experiences almost no force pulling it back toward more central regions of the data cloud. Its parameters become effectively “stuck” near the boundary, and the corresponding directions in the loss landscape remain flat enough to satisfy the BEoS condition. Our lower bound constructions exploit exactly this mechanism: by arranging many such trapped, boundary-supported neurons on disjoint caps, we obtain shattering networks that interpolate, remain dynamically stable, and yet are statistically hard to learn. Although our analysis is carried out for ReLU, where hard sparsity makes this effect particularly transparent, the underlying “weak-gradient trapping” mechanism suggests that similar phenomena may persist for other activations with rapidly decaying gradients away from their transition region.
How GD Adapts to Low-dimensionality Dynamically. Recall our notations: for a two-layer ReLU network
$f_{\bm{\theta}}({\bm{x}})=\sum_{k=1}^{K}v_{k}\,\phi({\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}-b_{k})+\beta$
trained on data $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ with empirical loss $\mathcal{L}({\bm{\theta}})=\frac{1}{2n}\sum_{i=1}^{n}\ell(f_{{\bm{\theta}}}({\bm{x}}_{i}),y_{i})$, the gradient with respect to a hidden weight ${\bm{w}}_{k}$ has the form
$\nabla_{{\bm{w}}_{k}}\mathcal{L}({\bm{\theta}})=\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}\bigl(f_{{\bm{\theta}}}({\bm{x}}_{i}),y_{i}\bigr)\,v_{k}\,\phi^{\prime}({\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}_{i}-b_{k})\,{\bm{x}}_{i}=\sum_{i=1}^{n}\alpha_{k,i}({\bm{\theta}})\,{\bm{x}}_{i}.$ (13)
Thus, at every step of gradient descent, the update of ${\bm{w}}_{k}$ is a linear combination of input vectors. In informal terms, the shape of the gradient field is inherited from the shape of the data cloud: the dynamics cannot move in arbitrary directions in parameter space, but only along directions induced by the training inputs and the neuron-specific activation pattern.
In the idealized case where all ${\bm{x}}_{i}$ lie in a linear subspace $V\subset\mathbb{R}^{d}$, (13) already suggests a form of intrinsic-dimension adaptation: the trajectory of each hidden weight lives in an affine translate of $V$, so any meaningful notion of effective complexity should depend on $\dim V$ rather than on the ambient dimension. However, as soon as we move to more realistic geometries—affine subspaces $a+V$ or mixtures of $m$-dimensional components—the global span of the data can easily be full-dimensional. In those settings, the observation that gradients lie in $\mathrm{span}\{{\bm{x}}_{i}\}$ is nearly vacuous: it no longer encodes the structured way in which the data constrain the dynamics, and by itself it does not yield intrinsic-dimension generalization bounds.
Our contribution is to make this stability-based picture interact explicitly with the geometry of the data. For structured distributions such as mixtures of $m$-dimensional balls, the data-dependent weight function $g_{\mathcal{D}}({\bm{u}},t)$ inherits a corresponding structure. A hyperplane that cuts through the thick interior of an $m$-dimensional component activates on many of its points. Neurons aligned with such hyperplanes receive large gradients and are strongly constrained by the stability condition, which is reflected in large values of $g_{\mathcal{D}}({\bm{u}},t)$. In contrast, hyperplanes that only touch shallow caps or skim the boundary fire on very few points; neurons aligned with these directions see much weaker “gradient pressure” and correspondingly small values of $g_{\mathcal{D}}({\bm{u}},t)$, allowing them to carry high norm and implement localized features.
In this way, the gradient-dynamics perspective can be summarized as
$\text{data geometry}\;\Longrightarrow\;\text{gradient geometry}\;\Longrightarrow\;\text{stability-induced regularization},$
with $g_{\mathcal{D}}$ acting as the bridge between dynamics and capacity control. Stability does not merely say that gradients live in the span of the data; through $g_{\mathcal{D}}$, it tells us which parts of the data geometry are expensive to fit and which parts can host high-norm, sparsely supported neurons. This refinement allows our analysis to turn the qualitative statement that “gradient trajectories are shaped by the data geometry” into quantitative, distribution-dependent generalization bounds that scale with the intrinsic dimension $m$ rather than the ambient dimension $d$.
@@ Appendix C Details of Experiments
(b) Sample-wise error versus half-space depth at epoch 5,0005{,}000
@@ D.2 Total Variation Semi-norm on Radon Domain
We now connect the (unweighted) variation semi-norm of shallow ReLU networks to an analytic description on the Radon domain. Our presentation follows [Parhi and Nowak, 2021, 2022, 2023, 2026, Parhi and Unser, 2024].
@@
$c_{d}\,f\;=\;\mathscr{R}^{*}\left\{\,\Lambda^{d-1}\mathscr{R}f\,\right\},$
where $\Lambda^{d-1}$ acts in the $t$-variable with Fourier symbol $\widehat{\Lambda^{d-1}\Phi}({\bm{u}},\omega)\propto|\omega|^{\,d-1}\,\hat{\Phi}({\bm{u}},\omega)$.
@@
where the fractional power is understood via the Fourier transform.
Proposition D.7 (Equivalence of seminorms on bounded domains [Parhi and Nowak, 2023]).
@@
Remark D.8.
For ReLU networks on bounded domains, the three viewpoints:
$\text{path-norm }\left\lVert f\right\rVert_{\mathrm{path}}\;\;\longleftrightarrow\;\;\text{unweighted variation }|f|_{\mathrm{V}}\;\;\longleftrightarrow\;\;\text{second-order Radon-TV }\mathscr{R}\mathrm{TV}^{2}(f)$
are equivalent up to affine functions.
@@ E.1 Function Space Viewpoint of Neural Networks Below the Edge of Stability
Remark E.3.
$\mathrm{V}_{g}$ is a weighted variation space in the sense of DeVore et al. [2025].
@@ E.2 Empirical Process for the Weight Function $g$
Taking the supremum over $({\bm{u}},t)\in\mathbb{S}^{d-1}\times[-1,1]$ and applying Lemmas E.5 and E.6 with a union bound yields
@@
Recall that for $Q\in\{\mathcal{D},\mathcal{P}\}$,
$g_{Q}({\bm{u}},t)=\min\{\tilde{g}_{Q}({\bm{u}},t),\tilde{g}_{Q}(-{\bm{u}},-t)\}.$
For any real numbers $a,b,c,d$,
$\big|\min\{a,b\}-\min\{c,d\}\big|\leq\max\{|a-c|,\;|b-d|\}.$
Hence
$\sup_{{\bm{u}}\in\mathbb{S}^{d-1},\,t\in[-1,1]}|g_{\mathcal{D}}({\bm{u}},t)-g_{\mathcal{P}}({\bm{u}},t)|\leq\sup_{{\bm{u}}\in\mathbb{S}^{d-1},\,t\in[-1,1]}|\tilde{g}_{\mathcal{D}}({\bm{u}},t)-\tilde{g}_{\mathcal{P}}({\bm{u}},t)|.$
Since $({\bm{u}},t)\mapsto(-{\bm{u}},-t)$ preserves $\mathbb{S}^{d-1}\times[-1,1]$, the previously established uniform deviation bound for $\tilde{g}_{\mathcal{D}}-\tilde{g}_{\mathcal{P}}$ applies to both terms inside the minimum, and therefore transfers to $g_{\mathcal{D}}-g_{\mathcal{P}}$ with the same rate and failure probability. ∎
@@ Appendix F Generalization Upper Bound: Mixture of Low-Dimensional Balls
In this section, we present the proof of Theorem 3.10. We formalize the mixture model by the following setting.
Assumption F.1 (Mixture of low-dimensional balls).
where $p_{j}>0$ are the mixture probabilities $\mathbb{P}({\bm{x}}\in V_{j})$ satisfying $\sum_{j=1}^{J}p_{j}=1$. Each component distribution $\mathcal{P}_{j}$ is the uniform distribution on the unit ball $\mathbb{B}_{1}^{V_{j}}:=\{{\bm{x}}\in V_{j}:\left\lVert{\bm{x}}\right\rVert_{2}\leq 1\}$. The corresponding labels $y$ are generated from a conditional distribution $\mathcal{P}(y|{\bm{x}})$ and are assumed to be bounded, i.e., $|y|\leq D$ for some constant $D>0$. Similarly, we define $\mathcal{P}_{j}({\bm{x}},y)=\mathcal{P}({\bm{x}},y\mid{\bm{x}}\in V_{j})$.
First, we prove the simple case of singe-subspace assumption ($J=1$) via Theorem F.3.
@@ F.1 Case: uniform distribution on unit disc of a linear subspace
By Lemma F.2, it remains to consider $\operatorname{proj}_{V}^{*}f_{{\bm{\theta}}}$. Let $\mathcal{D}\subset V$ and let $\bm{X}$} be supported on $V$. For any ${\bm{w}}\in\mathbb{R}^{d}$ with $\operatorname{proj}_{V}({\bm{w}})\neq 0$, set
${\bm{u}}:=\frac{{\bm{w}}}{\|{\bm{w}}\|},\qquad{\bm{u}}_{V}:=\frac{\operatorname{proj}_{V}({\bm{w}})}{\|\operatorname{proj}_{V}({\bm{w}})\|}.$
Since for all ${\bm{x}}\in V$,
${\bm{w}}^{\top}{\bm{x}}=\operatorname{proj}_{V}({\bm{w}})^{\top}{\bm{x}},$
we have
${\bm{w}}^{\top}{\bm{x}}>b\quad\Longleftrightarrow\quad\operatorname{proj}_{V}({\bm{w}})^{\top}{\bm{x}}>b.$
After normalization,
${\bm{u}}^{\top}{\bm{x}}>\frac{b}{\|{\bm{w}}\|}\quad\Longleftrightarrow\quad{\bm{u}}_{V}^{\top}{\bm{x}}>\frac{b}{\|\operatorname{proj}_{V}({\bm{w}})\|}.$
Hence
$g\!\left(\frac{{\bm{w}}}{\|{\bm{w}}\|},\frac{b}{\|{\bm{w}}\|}\right)=g\!\left(\frac{\operatorname{proj}_{V}({\bm{w}})}{\|\operatorname{proj}_{V}({\bm{w}})\|},\frac{b}{\|\operatorname{proj}_{V}({\bm{w}})\|}\right).$
Therefore, when the data are supported on $V$, the weight function depends only on directions in $V$. It suffices to consider the class
${\bm{\Theta}}^{V}_{g_{\mathcal{D}}}(\mathbb{B}_{1}^{V_{j}};M,C)=\left\{\operatorname{proj}_{V}^{*}f_{{\bm{\theta}}}\>\colon\>f_{{\bm{\theta}}}\in{\bm{\Theta}}_{g_{\mathcal{D}}}(\mathbb{B}_{1}^{V_{j}};M,C)\right\}.$
Therefore, we just need consider the case where the whole algorithm with any dataset sample from $V$ operates in $V$ and we get the result from [Liang et al., 2025, Theorem F.8] by replacing $\mathbb{R}^{d}$ with $V\cong\mathbb{R}^{m}$. ∎